OpenStudy (anonymous):

How would I use number properties to simplify the following expression. -5 + (5 + 3)

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

I think I need to use the additive inverse or additive identity.

OpenStudy (skullpatrol):

$$\Huge a+(b+c)=(a+b)+c$$$$\Huge -5+(5+3) = (-5+5)+3$$

OpenStudy (anonymous):

Is that the distributive property, and is that all I need to do?

OpenStudy (skullpatrol):

That is the associative property

OpenStudy (anonymous):

Oh

OpenStudy (skullpatrol):

$$\Huge -a+a=0$$ $$\Huge -5+5=0$$

OpenStudy (skullpatrol):

$$\Huge 0+a=a$$ $$\Huge 0 +3=3$$

OpenStudy (anonymous):

And I believe that's the additive inverse and the second one would be the additive identity

OpenStudy (skullpatrol):

a + (-a) = 0 is called the property of opposites a + 0 = a is called the identity property of addition

OpenStudy (anonymous):

Thanks!

OpenStudy (skullpatrol):

Thanks for asking :-)